RC meeting and Run Plan, Sunday December 11th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

(NO SUNDAY MEETING) RC Meeting Information

Current RC: Eric King [ericking (@jlab.org or @temple.edu)]
Time: Daily at 3:30PM (NO SUNDAY MEETING)
Location: Counting House Conference Room

Remote Connection Details

Shift Schedule

Link to Schedule

Until Shutdown

All shifts are full! :)

Accelerator Status

  • Accelerator was down at the end of OWL for about seventy minutes (75) at the end of OWL. [DT Event ID: 17471]
  • Accelerator was down at the end of SWING beginning of OWL for a total of two (2) hours. [DT Event ID: 17452]
  • Hall C valve/vacuum repair is complete; they're coming back online Tuesday; they will be priority hall next week.

Current Status

Taking production data on 3-He

Shift Summaries

Things That Happened

  • HH Hodoscope voltages dropped. Shift leader notified me around 5:30am. Resolved over phone. [4099867]
  • Scalar helicity graph from run 3191 showed all zeros. [4099876]
    • SL did not immediately notify me but resolved on own; unfortunately, the 50K replay was not looked at immediately.
    • SL had already proceeded with what I would have suggested and was told to notify me if the solution didn't work when beam was restored.

Beam Current

We've been happily running at 40uA.

  • I asked Sebastian to make a record of the ramp rates over the course of the past 8 hours during Saturday SWING and he went an extra step. [4099752]
    • Bill has taken this further and created a strip chart that tracks this continually that seems to work very nicely. [4099956}


Target polarization has been stable. List of the recent polarizations

Bill noticed that the wrong laser software was running on polhe1; he corrected the issue. [4100010]

Issues Resolved

  • Bill Henry has resolved the communication issue regarding the Moller HH magnet. Thank you, Bill. [4099942]
    • The remaining alarm handler issue is the unused GEMs. Bill spoke with Zeke about this and they agreed to bring the topic up in Monday's RC Meeting.

Issues remaining

  • BB GEM APV unreachable and cannot be repaired at this time.

Short Term Run Plan

  • Continue production run on He3 target, 40 uA
  • Use GEnII-noSBSGEMs DAQ config
  • Use ps3=0, ps5=6, others =-1
  • BBCal threshold at -366 mV, HCal threshold at -19 mV
  • NMR measurement every 3 - 4 hours, close to 4 hours is better

Planned Accesses to the Hall


Special Requests

Please email me ericking (@jlab.org or @temple.edu) if there's something to be added.

Notices in Effect for Shift Crew

  • Power down SBS, BB, and Detector HV before making an access to the Hall

Data Collection Status

  • 77 out of 141 days (~54.6%)
  • (NEEDS TO BE UPDATED) Total uncorrected accumulated charge on He3: 53.937 C out of 232C (23.25%)
    • 2-Pass: 13.467 C
    • 3-Pass: 40.470 C (@ end of SUNDAY DAY)